# MCMC Marginalisation Bias and $\Lambda$CDM tensions

### v2

@@ Abstract
Hubble tension is a problem in one-dimensional (1D) posteriors, since local $H_{0}$ determinations are only sensitive to a single parameter. Projected 1D posteriors for $\Lambda$CDM cosmological parameters become more non-Gaussian with increasing effective redshift when the model is fitted to redshift binned data in the late Universe. We explain mathematically why this non-Gaussianity arises and show using observational Hubble data (OHD) that Markov Chain Monte Carlo (MCMC) marginalisation leads to 1D posteriors that fail to track the $\chi^{2}$ minimum at $68\%$ confidence level in high redshift bins. To remedy this limitation of MCMC, we resort to profile distributions as a complementary technique. Doing so, we observe that $z\gtrsim 1$ cosmic chronometer (CC) data currently prefers a non-evolving (constant) Hubble parameter over a Planck-$\Lambda$CDM cosmology at $\sim 2\sigma$. Within the Hubble tension debate, it is imperative that subsamples of data sets with differing redshifts yield similar $H_{0}$ values. In addition, we confirm that MCMC degeneracies observed in 2D posteriors are not due to curves of constant $\chi^{2}$. Finally, on the assumption that the Planck-$\Lambda$CDM cosmological model is correct, using profile distributions we confirm a $>2\sigma$ discrepancy with Planck-$\Lambda$CDM in a combination of CC and baryon acoustic oscillations (BAO) data beyond $z\sim 1.5$. This confirms a discrepancy reported earlier with fresh methodology.
@@ 1 Introduction
The flat $\Lambda$CDM model is the minimal model that fits Cosmic Microwave Background (CMB) data. Remarkably, CMB data from the Planck satellite Planck:2018vyg constrains the $\Lambda$CDM model to sub-percent errors, thereby not only providing the strongest constraints, but also a prediction for cosmological probes in the late Universe. The unmitigated success of the $\Lambda$CDM model is that CMB, Type Ia supernovae (SN) Riess:1998cb ; Perlmutter:1998np and baryon acoustic oscillations (BAO) Eisenstein:2005su agree on a $\Lambda$CDM Universe that is approximately $30\%$ pressureless matter. Thus, one key prediction of the Planck-$\Lambda$CDM model agrees across early and late Universe cosmological probes. Given this non-trivial agreement, any discrepancies that arise elsewhere constitute puzzles.
Nevertheless, one cannot define any model for a dynamical system, especially a complicated system like the Universe, using data from a cosmic snapshot. [^1] Here, we mean CMB data with an effective redshift $z\sim 1100$. At best, one has a prediction and not a model. In recent years, key predictions of Planck data have been challenged by late Universe determinations of the Hubble constant $H_{0}$ Riess:2021jrx ; Freedman:2021ahq ; Pesce:2020xfe ; Blakeslee:2021rqi ; Kourkchi:2020iyz and the $S_{8}:=\sigma_{8}\sqrt{\Omega_{m}/0.3}$ parameter HSC:2018mrq ; KiDS:2020suj ; DES:2021wwk ; Boruah:2019icj ; Said:2020epb . Given the diversity of the late Universe probes (see reviews Perivolaropoulos:2021jda ; Abdalla:2022yfr ), it is highly unlikely that any single systematic can be found to explain the discrepancies. That being said, in astrophysics one can never preclude systematics; 3 decades after Phillips’ seminal paper Phillips:1993ng , we are still debating an ad hoc correction for the mass of the host galaxy in Type Ia SN NearbySupernovaFactory:2018qkd ; Kang:2019azh ; Brout:2020msh ; Lee:2021txi . Bearing in mind that Type Ia SN are one of our best understood cosmological probes, one quickly understands that any systematics debate may be endless.
@@
Within the context of $\Lambda$CDM tensions, it was recently observed that the integration constant from the Friedmann equations, aka the Hubble constant $H_{0}$, picks up redshift dependence whenever our model assumption - required to close the Friedmann equations - disagrees with the Hubble parameter $H(z)$ extracted from observations Krishnan:2020vaf ; Krishnan:2022fzz . Similarly, $\rho_{m0}=H_{0}^{2}\Omega_{m}$, an integration constant of the matter continuity equation, implies matter density $\Omega_{m}$ is a mathematically constant quantity. These are irrefutable predictions from mathematics, i. e. a prediction that is robust to systematics. However, observationally $H_{0}$ and $\Omega_{m}$ are model fitting parameters and nothing precludes them picking up redshift dependence (except of course if one assumes they do not!), and providing a signature of model breakdown. If this happens in the late Universe within the $\Lambda$CDM model, $H_{0}$ is correlated with matter density $\Omega_{m}$, while $\Omega_{m}$ is correlated with $S_{8}\propto\sigma_{8}\sqrt{\Omega_{m}}$. Thus, there is at least one simple scenario, namely redshift evolution of cosmological parameters in the late Universe, where “$H_{0}$ tension” and “$S_{8}$ tension” are not independent and simply symptoms of $\Lambda$CDM model breakdown.
@@
The next relevant question is, where is the evidence for evolving cosmological parameters in the late Universe? Starting with strong lensing time delay Wong:2019kwg ; Millon:2019slk , [^2] Systematics are explored in Millon:2019slk and the descending trend is not an obvious systematic. The lensed system RXJ1131-1231 Sluse:2003iy , which partly drives the trend, has recently been re-analysed using spatially resolved stellar kinematics of the host galaxy Shajib:2023uig , and the higher $H_{0}$ value remains robust, admittedly with inflated errors. As TDCOSMO project to analyse 40 lenses, the prospect of a discovery of a descending $H_{0}$ trend assuming the $\Lambda$CDM model remain strong. descending trends of $H_{0}$ with redshift have been reported in Type Ia SN Dainotti:2021pqg ; Colgain:2022nlb ; Colgain:2022rxy ; Malekjani:2023dky ; Hu:2022kes ; Jia:2022ycc and combinations of data sets Krishnan:2020obg ; Dainotti:2022bzg . On the other hand, larger values of $\Omega_{m}$ have been noted in high redshift observables, primarily quasars (QSOs) Risaliti:2015zla ; Risaliti:2018reu ; Lusso:2020pdb ; Yang:2019vgk ; Khadka:2020vlh ; Khadka:2020tlm ; Khadka:2021xcc ; Pourojaghi:2022zrh , [^3] Just as with Type Ia SN, the systematics of QSOs are being investigated Zajacek:2023qjm . but also Type Ia SN Colgain:2022nlb ; Colgain:2022rxy ; Malekjani:2023dky ; Pasten:2023rpc (see also Wagner:2022etu ; Sakr:2023hrl ). Note, as emphasised earlier, if $H_{0}$ evolves at the background level, correlated fitting parameters are expected to also evolve. Moreover, mock analysis within the $\Lambda$CDM setting reveals that evolution of best fit $(H_{0},\Omega_{m})$ parameters cannot be precluded, and conversely possesses a finite likelihood, in either observational Hubble data (OHD) or angular diameter distance data or luminosity distance data Colgain:2022tql . We stress that this result rests on mock analysis; it represents a purely mathematical statement about the $\Lambda$CDM model that is independent of systematics.
@@
Separately, at the perturbative level, redshift evolution of $S_{8}$ or $\sigma_{8}$ has been reported in galaxy cluster number counts and Lyman-$\alpha$ spectra Esposito:2022plo , $f\sigma_{8}$ constraints from peculiar velocities and redshift space distortions (RSD) Adil:2023jtu , comparison between weak HSC:2018mrq ; KiDS:2020suj ; DES:2021wwk and CMB lensing ACT:2023dou ; ACT:2023kun . What is important here is that these observations appear to restrict the evolution in $S_{8}$ to the late Universe. In ACT:2023ipp the possibility was raised that “tracers at higher redshift and probing larger scales prefer higher $S_{8}$. [^4] There are also conflicting observations of high redshift $\sigma_{8}$ or $S_{8}$ values that are lower than Planck in the late Universe Miyatake:2021qjr ; Alonso:2023guh , so either this trend is not universal, or systematics are at play. Nevertheless, one can argue against evolution with scale on the grounds that cosmic shear HSC:2018mrq ; KiDS:2020suj ; DES:2021wwk , which is sensitive to smaller scales (larger $k$), and peculiar velocity constraints Boruah:2019icj ; Said:2020epb , which are sensitive to larger scales (smaller $k$), both prefer lower values of $S_{8}$. Moreover, both galaxy clusters and Lyman-$\alpha$ spectra are expected to probe similar scales. [^5] We thank Matteo Viel for correspondence on this point. Thus, if systematics are not impacting results, then redshift evolution is the only point of agreement in the observations Esposito:2022plo ; Adil:2023jtu ; HSC:2018mrq ; KiDS:2020suj ; DES:2021wwk ; ACT:2023dou ; ACT:2023kun ; ACT:2023ipp . Note also that redshift is more fundamental than scale in FLRW cosmology; one must solve the Friedmann equations in either time or redshift before one contemplates any discussion of scale.
@@
The purpose of this letter is to revisit the analysis presented in Colgain:2022rxy ; Colgain:2022tql , where the evidence for evolution was quantified on the basis of mock simulations and not Markov Chain Monte Carlo (MCMC), the technique most familiar in cosmology. The fundamental problem is that once one bins low redshift data and studies evolution of cosmological parameters with bin redshift, one quickly encounters degeneracies and projection effects in MCMC analyses. These effects are not just the preserve of exotic models Herold:2021ksg ; Gomez-Valent:2022hkb ; Meiers:2023gft , such as Early Dark Energy (EDE) Poulin:2018cxd ; Niedermann:2019olb , and happen in the simplest model when one bins data. The most striking demonstration of the resulting bias is that the peaks of MCMC posteriors no longer coincide with the minimum of the likelihood (see Gomez-Valent:2022hkb ). Ultimately, this bias is expected because one is working in a regime of the $\Lambda$CDM model with projected 1D non-Gaussian probability distributions Colgain:2022tql (see also Colgain:2022rxy ). We remind the reader once again that Hubble tension is a problem in 1D posteriors, so projection cannot be avoided if the goal is to compare cosmological and local $H_{0}$ constraints.
@@
The structure of this paper is as follows. In section 2 we confirm the bias in MCMC marginalisation. In section 3 we introduce profile distributions (PDs) Gomez-Valent:2022hkb as a means of addressing the bias and confirm that the statistical significance of discrepancies from mock simulations agree well with PD analysis. In section 4 we study whether a degeneracy in MCMC posteriors in the Bayesian approach translates into a constant $\chi^{2}$ curve in the Frequentist approach. In section 5, we revisit and confirm the high redshift OHD tensions reported in Colgain:2022rxy . We end in section 6 with concluding remarks.
2 A bias in MCMC marginalisation
In this section we illustrate a bias in MCMC marginalisation that arises in the (flat) $\Lambda$CDM model when data is binned by redshift. This bias is inherent to the $\Lambda$CDM model and arises due to a projection effect of banana-shaped contours when one confronts the model to either high redshift angular diameter (alternatively luminosity) distance constraints or OHD Colgain:2022rxy ; Colgain:2022tql . Given its generality, it is not surprising that it also impacts MCMC inferences.
@@ 2.1 Mathematical Arguments
More concretely, taking a large $z$ limit we see that high redshift data can only constrain the combination $\rho_{m0}=H_{0}^{2}{\Omega_{m}}$. For data in a redshift bin with large $z$, $z\gg 0$, this means that one can only constrain the combination $\Omega_{m}h^{2}$ ($h:=H_{0}/100)$, but $H_{0}$ and $\Omega_{m}$ remain largely unconstrained. Alternatively put, for any given $\Omega_{m}h^{2}$ constraint, there is a large number of permissible $(H_{0},\Omega_{m})$ pairs. Translated into a probability density function (PDF), this is simply the statement that in a very high redshift bin with $z\gg 0$, one expects almost uniform/flat distributions for $H_{0}$ and $\Omega_{m}$ with the model (1).
@@
Here what is important to us is the observation that these PDFs must flatten in a non-Gaussian manner. To appreciate this fact, we observe that high redshift OHD only constrains $\Omega_{m}h^{2}$ well. [^6] Note that observables like SN or QSO that measure $D_{L}(z)=c(1+z)\int_{0}^{z}\textrm{d}z^{\prime}/H(z^{\prime})$ are mainly sensitive to the low redshift part of $H(z)$, i. e. the combination $H_{0}^{2}(1-\Omega_{m})$, and in this sense they are complementary to the OHD data which is more sensitive to high redshift part of $H(z)$, $H_{0}^{2}\Omega_{m}$. The complementarity can be demonstrated by combining $H(z)$ and $D_{L}(z)$ constraints and checking that one recovers mock data input parameters in all redshift bins Colgain:2022tql . For this reason, best fit parameters are constrained to a $\Omega_{m}h^{2}=\textrm{constant}$ curve in the $(H_{0},\Omega_{m})$-plane. The almost flat $H_{0}$ and $\Omega_{m}$ PDFs can only arise if this curve stretches in the $(H_{0},\Omega_{m})$-plane. As a result of this stretching, one ends up with a relatively uniform distribution on a curve. At the extremes of the curve, one finds a distribution of large $H_{0}$ values, which do not differ greatly in $\Omega_{m}$, and they get projected to a peak at small values on the $\Omega_{m}$ axis. Conversely, at the other end of the curve, one finds a distribution of small $\Omega_{m}$ values, which do not differ greatly in $H_{0}$, and they get projected onto a peak at large values on the $H_{0}$ axis. This is a “projection effect” in common cosmology parlance. It is driven by the irrelevance of the DE sector at high redshift and the constraint $\Omega_{m}h^{2}=\textrm{constant}$ from the $\Lambda$CDM model (1). Together these features distort projected 1D distributions away from a Gaussian configuration.
@@
Thus, simply by binning and fitting OHD to the $\Lambda$CDM model one enters a non-Gaussian regime as the effective redshift of the bin increases. This effect, which is expected from the purely mathematical arguments above, has been confirmed in mock data Colgain:2022rxy ; Colgain:2022tql , and in line with expectations, we demonstrate that it impacts MCMC inferences with observed data in the next subsection.

### v3

# A comparison of Bayesian and frequentist confidence intervals in the presence of a late Universe degeneracy

@@ Abstract
Hubble tension is a problem in one-dimensional (1D) posteriors, since local $H_{0}$ determinations are only sensitive to a single parameter. Projected 1D posteriors for $\Lambda$CDM cosmological parameters become more non-Gaussian with increasing effective redshift when the model is fitted to redshift-binned data in the late Universe. We explain mathematically why this non-Gaussianity arises and show using observational Hubble data (OHD) that Markov Chain Monte Carlo (MCMC) marginalisation leads to 1D posteriors that fail to track the $\chi^{2}$ minimum at $68\%$ confidence level in high redshift bins. To gain a second perspective, we resort to profile likelihoods as a complementary technique. Doing so, we observe that $z\gtrsim 1$ cosmic chronometer (CC) data currently prefers a non-evolving (constant) Hubble parameter over a Planck-$\Lambda$CDM cosmology at $\sim 2\sigma$. Within the Hubble tension debate, it is imperative that subsamples of data sets with differing redshifts yield similar $H_{0}$ values. In addition, we confirm that MCMC degeneracies observed in 2D posteriors are not due to curves of constant $\chi^{2}$. Finally, on the assumption that the Planck-$\Lambda$CDM cosmological model is correct, using profile likelihoods we confirm a $>2\sigma$ discrepancy with Planck-$\Lambda$CDM in a combination of CC and baryon acoustic oscillations (BAO) data beyond $z\sim 1.5$. This confirms a discrepancy reported earlier with fresh methodology.
@@ 1 Introduction
The flat $\Lambda$CDM model is the minimal model that fits Cosmic Microwave Background (CMB) data. Remarkably, CMB data from the Planck satellite Planck:2018vyg constrains the $\Lambda$CDM model to sub-percent errors, thereby not only providing the strongest constraints, but also a prediction for cosmological probes in the late Universe. The unmitigated success of the $\Lambda$CDM model is that CMB, Type Ia supernovae (SNe) Riess:1998cb ; Perlmutter:1998np and baryon acoustic oscillations (BAO) Eisenstein:2005su agree on a $\Lambda$CDM Universe that is approximately $30\%$ pressureless matter. Thus, one key prediction of the Planck-$\Lambda$CDM model agrees across early and late Universe cosmological probes. Given this non-trivial agreement, any discrepancies that arise elsewhere constitute puzzles.
Nevertheless, one cannot define any model for a dynamical system, especially a complicated system like the Universe, using data from a cosmic snapshot. [^1] Here, we mean CMB data with an effective redshift $z\sim 1100$. At best, one has a prediction and not a model. In recent years, key predictions of Planck data have been challenged by late Universe determinations of the Hubble constant $H_{0}$ Riess:2021jrx ; Freedman:2021ahq ; Pesce:2020xfe ; Blakeslee:2021rqi ; Kourkchi:2020iyz and the $S_{8}:=\sigma_{8}\sqrt{\Omega_{m}/0.3}$ parameter HSC:2018mrq ; KiDS:2020suj ; DES:2021wwk ; Boruah:2019icj ; Said:2020epb . Given the diversity of the late Universe probes (see reviews Perivolaropoulos:2021jda ; Abdalla:2022yfr ), it is highly unlikely that any single systematic can be found to explain the discrepancies. That being said, in astrophysics one can never preclude systematics; 3 decades after Phillips’ seminal paper Phillips:1993ng , we are still debating an ad hoc correction for the mass of the host galaxy in Type Ia SNe NearbySupernovaFactory:2018qkd ; Kang:2019azh ; Brout:2020msh ; Lee:2021txi . Bearing in mind that Type Ia SNe are one of our best understood cosmological probes, one quickly understands that any systematics debate may be endless.
@@
Within the context of the $\Lambda$CDM model, the Hubble constant $H_{0}$ arises mathematically as an integration constant when one solves the Friedmann equations. This provides a mathematical perspective on $H_{0}$ Krishnan:2020vaf ; Krishnan:2022fzz . Nevertheless, there is a second observational perspective; $H_{0}$ is a fitting parameter in the $\Lambda$CDM model. As a result, one can perform a consistency check of the $\Lambda$CDM model by confronting it to observational data in different redshift ranges or cosmological epochs Krishnan:2020vaf ; Krishnan:2022fzz ; Colgain:2022nlb ; Colgain:2022rxy . If one recovers the same $H_{0}$ value at all redshifts, then mathematics and observation are in sync. In contrast, if one does not, the data and the $\Lambda$CDM model are no longer in agreement. Similarly, $\rho_{m0}=H_{0}^{2}\Omega_{m}$, an integration constant of the matter continuity equation, implies matter density $\Omega_{m}$ is a mathematically constant quantity, so the same logic applies. In the late Universe within the $\Lambda$CDM model, $H_{0}$ is correlated with matter density $\Omega_{m}$, while $\Omega_{m}$ is correlated with $S_{8}\propto\sigma_{8}\sqrt{\Omega_{m}}$. Thus, there is at least one simple scenario, namely redshift evolution of cosmological fitting parameters in the late Universe, where “$H_{0}$ tension” and “$S_{8}$ tension” are not independent and simply symptoms of $\Lambda$CDM model breakdown Colgain:2022nlb ; Colgain:2022rxy ; Colgain:2022tql .
@@
The next relevant question is, where is the evidence for evolving cosmological fitting parameters in the late Universe? Starting with strong lensing time delay Wong:2019kwg ; Millon:2019slk , [^2] Systematics are explored in Millon:2019slk and the descending trend is not an obvious systematic. The lensed system RXJ1131-1231 Sluse:2003iy , which partly drives the trend, has recently been re-analysed using spatially resolved stellar kinematics of the host galaxy Shajib:2023uig , and the higher $H_{0}$ value remains robust, admittedly with inflated errors. As TDCOSMO project to analyse 40 lenses, the prospect of a discovery of a descending $H_{0}$ trend assuming the $\Lambda$CDM model remain strong. descending trends of $H_{0}$ with redshift have been reported in Type Ia SNe Dainotti:2021pqg ; Colgain:2022nlb ; Colgain:2022rxy ; Malekjani:2023dky ; Hu:2022kes ; Jia:2022ycc and combinations of data sets Krishnan:2020obg ; Dainotti:2022bzg . On the other hand, larger values of $\Omega_{m}$ have been noted in high redshift observables, primarily quasars (QSOs) Risaliti:2015zla ; Risaliti:2018reu ; Lusso:2020pdb ; Yang:2019vgk ; Khadka:2020vlh ; Khadka:2020tlm ; Khadka:2021xcc ; Pourojaghi:2022zrh , [^3] Just as with Type Ia SNe, the systematics of QSOs are being investigated Zajacek:2023qjm . but also Type Ia SNe Colgain:2022nlb ; Colgain:2022rxy ; Malekjani:2023dky ; Pasten:2023rpc (see also Wagner:2022etu ; Sakr:2023hrl ). Note, as emphasised earlier, if $H_{0}$ as a fitting parameter evolves at the background level, which is only possible if the $\Lambda$CDM model has broken down, correlated fitting parameters are expected to also evolve. Moreover, mock analysis within the $\Lambda$CDM setting reveals that evolution of best fit $(H_{0},\Omega_{m})$ parameters cannot be precluded, and conversely possesses a finite likelihood, in either observational Hubble data (OHD) or angular diameter distance data or luminosity distance data Colgain:2022tql . We stress that this result rests on mock analysis; it represents a purely mathematical statement about the $\Lambda$CDM model that is independent of systematics.
@@
Separately, at the perturbative level, redshift evolution of $S_{8}$ or $\sigma_{8}$ has been reported in galaxy cluster number counts and Lyman-$\alpha$ spectra Esposito:2022plo , $f\sigma_{8}$ constraints from peculiar velocities and redshift space distortions (RSD) Adil:2023jtu , comparison between weak HSC:2018mrq ; KiDS:2020suj ; DES:2021wwk and CMB lensing ACT:2023dou ; ACT:2023kun (see also Tutusaus:2023aux ). What is important here is that these observations appear to restrict the evolution in $S_{8}$ to the late Universe. In ACT:2023ipp the possibility was raised that “tracers at higher redshift and probing larger scales prefer higher $S_{8}$. [^4] There are also conflicting observations of high redshift $\sigma_{8}$ or $S_{8}$ values that are lower than Planck in the late Universe Miyatake:2021qjr ; Alonso:2023guh , so either this trend is not universal, or systematics are at play. Nevertheless, one can argue against evolution with scale on the grounds that cosmic shear HSC:2018mrq ; KiDS:2020suj ; DES:2021wwk , which is sensitive to smaller scales (larger $k$), and peculiar velocity constraints Boruah:2019icj ; Said:2020epb , which are sensitive to larger scales (smaller $k$), both prefer lower values of $S_{8}$. Moreover, both galaxy clusters and Lyman-$\alpha$ spectra are expected to probe similar scales. [^5] We thank Matteo Viel for correspondence on this point. Thus, if systematics are not impacting results, then redshift evolution is the only point of agreement in the observations Esposito:2022plo ; Adil:2023jtu ; HSC:2018mrq ; KiDS:2020suj ; DES:2021wwk ; ACT:2023dou ; ACT:2023kun ; ACT:2023ipp . Note also that redshift is more fundamental than scale in FLRW cosmology; one must solve the Friedmann equations in either time or redshift before one contemplates any discussion of scale.
@@
The purpose of this letter is to revisit the analysis presented in Colgain:2022rxy ; Colgain:2022tql , where the evidence for evolution in fitting parameters was quantified on the basis of mock simulations and not Markov Chain Monte Carlo (MCMC), the technique most familiar in cosmology. The fundamental problem is that once one bins low redshift data and studies evolution of cosmological fitting parameters with bin redshift, one quickly encounters degeneracies and projection effects in MCMC analyses. These effects are not just the preserve of exotic models Herold:2021ksg ; Gomez-Valent:2022hkb ; Meiers:2023gft , such as Early Dark Energy (EDE) Poulin:2018cxd ; Niedermann:2019olb , and happen in the simplest model when one bins data. The most striking demonstration of the effect is that the peaks of MCMC posteriors no longer coincide with the minimum of the likelihood (see Gomez-Valent:2022hkb ). Ultimately, this outcome may be expected because one is working in a regime of the $\Lambda$CDM model with projected 1D non-Gaussian probability distributions Colgain:2022tql (see also Colgain:2022rxy ). Note, this is not a problem with MCMC marginalisation per se, but a deeper problem that impacts MCMC marginalisation. We remind the reader once again that Hubble tension is a problem in 1D posteriors, so projection cannot be avoided if the goal is to compare cosmological and local $H_{0}$ constraints.
@@
The structure of this paper is as follows. In section 2 we identify the problem prompted by a late Universe degeneracy in the $\Lambda$CDM model. In section 3 we introduce profile distributions (PDs) Gomez-Valent:2022hkb , a mild variant of frequentist profile likelihoods Trotta:2017wnx , as a means of addressing the effect and confirm that the statistical significance of discrepancies from mock simulations agree well with PD analysis. In section 4 we study whether a degeneracy in MCMC posteriors in the Bayesian approach translates into a constant $\chi^{2}$ curve in the frequentist approach. In section 5, we revisit and confirm the high redshift OHD tensions reported in Colgain:2022rxy . We end in section 6 with concluding remarks.
2 A late Universe degeneracy
In this section we illustrate a problem that arises in the (flat) $\Lambda$CDM model when OHD is binned by redshift. Our treatment through to the end of section 2.2 overlaps with Colgain:2022tql , where exclusively mock data was considered. Readers familiar with the observations in Colgain:2022tql may skip directly to subsection 2.3. Subsection 2.3 onwards discusses features in observed cosmic chronometer (CC) data. Our analysis in subsection 2.3 onwards overlaps with Colgain:2022rxy but is distinct due to the absence of external priors and BAO data. We reintroduce priors/BAO later in section 5 to comment on a discrepancy with the Planck-$\Lambda$CDM cosmology and how the results compare with Colgain:2022rxy .
Returning to our main focus, the basic problem is that as one fits $H(z)$ constraints in bins of increasing effective redshift, the late Universe $\Lambda$CDM Hubble parameter $H(z)$ transitions from a model with two degrees of freedom $(H_{0},\Omega_{m})$ to a model where only a single degree of freedom $\Omega_{m}h^{2}$ ($h:=H_{0}/100$) is well constrained. In the cosmology literature, this is presented as the idea that “high redshift data can only constrain $\Omega_{m}h^{2}$”. This statement is supported by banana-shaped contours in $(H_{0},\Omega_{m})$ MCMC posteriors. Throughout, little consideration has been given to whether Bayesian and frequentist methods agree on the “high redshifts” where the late Universe $\Lambda$CDM model loses the additional degree of freedom. Noting that the transition is inevitable, but gradual, the point of this paper is to explore whether Bayesian and frequentist methods lose sensitivity to the full parameter space at the same rate or in step. We will demonstrate that this is not the case and that frequentist methods can identify constraints in redshift ranges where MCMC posteriors are unconstrained and thus inconclusive. Alternatively put, within our assumptions and setup, we demonstrate the frequentist methods lead to better constrained confidence intervals than Bayesian credible intervals when a degeneracy is present. Moreover, unconstrained MCMC posteriors manifest themselves as projection effects in marginalised 1D posteriors, which biases MCMC inferences. Separately, we confirm that in settings where the 2D parameter space is well constrained, as is evident by Gaussian MCMC posteriors, that Bayesian and frequentist methods agree.
@@ 2.1 Mathematical Arguments
More concretely, taking a large $z$ limit we see that high redshift data can only constrain the combination $\rho_{m0}=H_{0}^{2}{\Omega_{m}}$. For data in a redshift bin with large $z$, $z\gg 0$, this means that one can only constrain the combination $\Omega_{m}h^{2}$, but $H_{0}$ and $\Omega_{m}$ remain largely unconstrained. Alternatively put, for any given $\Omega_{m}h^{2}$ constraint, there is a large number of permissible $(H_{0},\Omega_{m})$ pairs. Translated into a probability density function (PDF), this is simply the statement that in a very high redshift bin with $z\gg 0$, one expects almost uniform/flat distributions for $H_{0}$ and $\Omega_{m}$ with the model (1).
@@
Here what is important to us is the observation that these PDFs must flatten in a non-Gaussian manner. To appreciate this fact, we observe that high redshift OHD only constrains $\Omega_{m}h^{2}$ well. [^6] Note that observables like SNe or QSOs that measure $D_{L}(z)=c(1+z)\int_{0}^{z}\textrm{d}z^{\prime}/H(z^{\prime})$ are mainly sensitive to the low redshift part of $H(z)$, i. e. the combination $H_{0}^{2}(1-\Omega_{m})$, and in this sense they are complementary to the OHD data which is more sensitive to high redshift part of $H(z)$, $H_{0}^{2}\Omega_{m}$. The complementarity can be demonstrated by combining $H(z)$ and $D_{L}(z)$ constraints and checking that one recovers mock data input parameters in all redshift bins Colgain:2022tql . For this reason, best fit parameters are constrained to a $\Omega_{m}h^{2}=\textrm{constant}$ curve in the $(H_{0},\Omega_{m})$-plane. The almost flat $H_{0}$ and $\Omega_{m}$ PDFs can only arise if this curve stretches in the $(H_{0},\Omega_{m})$-plane. As a result of this stretching, one ends up with a relatively uniform distribution on a curve. At the extremes of the curve, one finds a distribution of large $H_{0}$ values, which do not differ greatly in $\Omega_{m}$, and they get projected to a peak at small values on the $\Omega_{m}$ axis. Conversely, at the other end of the curve, one finds a distribution of small $\Omega_{m}$ values, which do not differ greatly in $H_{0}$, and they get projected onto a peak at large values on the $H_{0}$ axis. This is a “projection effect” in common cosmology parlance. It is driven by the irrelevance of the DE sector at high redshift and the constraint $\Omega_{m}h^{2}=\textrm{constant}$ from the $\Lambda$CDM model (1). Together these features distort projected 1D distributions away from a Gaussian configuration.
@@
Thus, simply by binning and fitting OHD to the $\Lambda$CDM model one enters a non-Gaussian regime as the effective redshift of the bin increases. As explained, this non-Gaussianity can be traced to the $\Lambda$CDM model transitioning from a 2D model at lower redshifts to an effective 1D model at higher redshifts. This effect, which is expected from the purely mathematical arguments above, has been confirmed in mock data Colgain:2022rxy ; Colgain:2022tql , and in line with expectations, we demonstrate how it impacts MCMC inferences with observed data in the next subsection.